# Learning from Linear Algebra: A Graph Neural Network Approach to Preconditioner Design for

### v2

@@ 5.1 Preconditioners comparison
In our experiments, the value of the correction coefficient $\alpha$ is always negative when training with loss (3). Interestingly, when we train PreCorrector with loss (2), $\alpha$ is always positive. Detailed values of $\alpha$ can be found in the Appendix A.8.
Finally, certain engineering problems require iterative solvers to be run multiple times on the same matrices. For these problems, the preconditioner’s precomputation overhead contributes less to the total time. We also observe a good generalization of our approach when transferring our preconditioner between grids and datasets (see Appendix A.4). The transferability of the PreCorrector allows both to use the trained PreCorrector on similar problems and to train the PreCorrector on the easier problems with inference on the hard ones.
Table 3: Comparison on Poisson equation: classical algorithms for IC(00), GNN from (Li et al., 2023) and PreCorrector. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{IC}(0)$ $1.9\cdot 10^{-4}$ $0.134\pm 0.001$ $0.166\pm 0.001$ $0.199\pm 0.001$
($62\pm 0.0$) ($77\pm 0.0$) ($93\pm 0.0$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.093}\pm 0.072$ $\textbf{0.118}\pm 0.072$ $\textbf{0.142}\pm 0.072$
($31\pm 0.3$) ($41\pm 0.4$) ($50\pm 0.1$)
(Li et al., 2023) $2.5\cdot 10^{-3}$ $\underline{0.106}\pm 0.076$ $\underline{0.131}\pm 0.077$ $\underline{0.154}\pm 0.077$
($33\pm 0.0$) ($42\pm 0.0$) ($50\pm 0.0$)
$128\times 128$ $\text{IC}(0)$ $4.6\cdot 10^{-4}$ $0.901\pm 0.116$ $1.223\pm 0.156$ $1.412\pm 0.181$
($115\pm 0.0$) ($157\pm 0.5$) ($181\pm 0.0$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.391}\pm 0.085$ $\textbf{0.497}\pm 0.093$ $\textbf{0.594}\pm 0.100$
($48\pm 0.7$) ($61\pm 0.2$) ($73\pm 0.5$)
(Li et al., 2023) $2.5\cdot 10^{-3}$ $\underline{0.430}\pm 0.078$ $\underline{0.537}\pm 0.079$ $\underline{0.645}\pm 0.081$
($50\pm 0.0$) ($63\pm 0.0$) ($76\pm 0.0$)
Table 4: Comparison on diffusion equation with variance $0.1$: classical algorithms for IC(00), GNN from (Li et al., 2023) and PreCorrector. 1)1)1 ) None of the test linear systems converged to $10^{-3}$ tolerance in $300$ iterations.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$32\times 32$ $\text{IC}(0)$ $7.2\cdot 10^{-5}$ $\underline{0.043}\pm 0.001$ $\underline{0.055}\pm 0.002$ $\underline{0.066}\pm 0.002$
($32\pm 0.6$) ($42\pm 0.7$) ($50\pm 0.6$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $1.6\cdot 10^{-3}$ $\textbf{0.040}\pm 0.092$ $\textbf{0.050}\pm 0.092$ $\textbf{0.060}\pm 0.092$
($21\pm 0.4$) ($28\pm 0.4$) ($35\pm 0.4$)
(Li et al., 2023) $2.5\cdot 10^{-3}$ $0.088\pm 0.070$ $0.113\pm 0.071$ $0.135\pm 0.071$
($60\pm 5.9$) ($79\pm 7.6$) ($96\pm 9.1$)
$64\times 64$ $\text{IC}(0)$ $1.6\cdot 10^{-4}$ $0.145\pm 0.010$ $0.185\pm 0.012$ $0.218\pm 0.014$
($65\pm 1.4$) ($84\pm 1.0$) ($99\pm 0.9$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.099}\pm 0.087$ $\textbf{0.126}\pm 0.088$ $\textbf{0.151}\pm 0.089$
($33\pm 0.6$) ($43\pm 0.6$) ($53\pm 0.6$)
(Li et al., 2023) $2.5\cdot 10^{-3}$ nan1 nan1 nan1
@@
Results for GNN from (Li et al., 2023) are obtained with the same GNN hyperparameters and architecture as reported in the original paper with only one change: GNN is initialized with all weights and biases equal to zero and trained untill convergence. Without this change in architecture, the training did not converge. We also observed unstable training of the GNN from (Li et al., 2023) and unpredictable quality of the preconditioner depending on the training length. In our experiments, we could get a better preconditioner if we did not train until convergence. However, we could not predict the time for early stopping except by directly computing $\kappa(P^{-1}A)$, which is too expensive.
5.2 Loss function
As mentioned in Section 2, one should focus on approximating the low frequency components. The loss (3) does indeed reduce the distance between the extreme eigenvalues compared to IC(0) which can be observed in Table 5. Moreover, the gap between the extreme eigenvalues is covered by the increase in the minimum eigenvalue, which supports the hypothesis of low frequency cancellation. The maximum eigenvalue also increase, but by a much smaller order of magnitude.
In Table 5 we have also included recently obtained bounds for the minimum and maximum eigenvalues (Häusner et al., 2024):
$\begin{split}\lambda_{\text{min}}&\geq\|PA^{-1}\|_{F}^{-1}\geq\big{(}\|PA^{-1}-I\|_{F}+1\big{)}^{-1}\\ \lambda_{\text{max}}&\leq\|A-P\|_{2}\|P^{-1}\|_{2}+1\end{split}\,\,\,.$ (10)
At the same time, the preconditioner trained with the loss function (2) without weighting with $A^{-1}$ gives a worse effect on spectrum of $A$ (Table 6) and takes much more time to converge (Figure 2). Different distributions of the eigenvalues can be found in Figure 3 in the Appendix A.3
@@ 5.2 Loss function
[…]
@@ Appendix A Appendix
A.1 Additional experiments with IC(0) and ICt(1) preconditioners
Table 7: Comparison on diffusion equation with variance $0.1$: classical algorithms for IC(00) and ICt($1$) and PreCorrector. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{IC}(0)$ $1.6\cdot 10^{-4}$ $0.145\pm 0.010$ $0.185\pm 0.012$ $0.218\pm 0.014$
($65\pm 1.4$) ($84\pm 1.0$) ($99\pm 0.9$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.099}\pm 0.087$ $\textbf{0.126}\pm 0.088$ $\textbf{0.151}\pm 0.089$
($33\pm 0.6$) ($43\pm 0.6$) ($53\pm 0.6$)
$\text{ICt}(1)$ $8.6\cdot 10^{-4}$ $0.090\pm 0.0020$ $0.115\pm 0.002$ $0.136\pm 0.001$
($40\pm 0.9$) ($51\pm 0.8$) ($61\pm 0.6$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $2.3\cdot 10^{-3}$ $\textbf{0.088}\pm 0.094$ $\textbf{0.112}\pm 0.094$ $\textbf{0.134}\pm 0.094$
($23\pm 0.5$) ($31\pm 0.4$) ($38\pm 0.4$)
$128\times 128$ $\text{IC}(0)$ $4.6\cdot 10^{-4}$ $0.992\pm 0.153$ $1.267\pm 0.191$ $1.489\pm 0.221$
($131\pm 3.2$) ($168\pm 1.9$) ($198\pm 1.9$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.425}\pm 0.080$ $\textbf{0.538}\pm 0.086$ $\textbf{0.643}\pm 0.094$
($52\pm 1.5$) ($66\pm 1.3$) ($80\pm 1.6$)
$\text{ICt}(1)$ $3.4\cdot 10^{-3}$ $0.555\pm 0.074$ $0.709\pm 0.092$ $0.833\pm 0.109$
($80\pm 1.9$) ($102\pm 1.3$) ($120\pm 1.1$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $5.0\cdot 10^{-3}$ $\textbf{0.367}\pm 0.069$ $\textbf{0.469}\pm 0.071$ $\textbf{0.560}\pm 0.074$
($37\pm 0.9$) ($48\pm 0.8$) ($58\pm 0.8$)
Table 8: Comparison on diffusion equation with variance $0.5$: classical algorithms for IC(00) and ICt($1$) and PreCorrector. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{IC}(0)$ $1.6\cdot 10^{-4}$ $0.165\pm 0.067$ $0.207\pm 0.069$ $0.242\pm 0.070$
($73\pm 2.1$) ($93\pm 2.1$) ($109\pm 2.2$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.098}\pm 0.095$ $\textbf{0.124}\pm 0.096$ $\textbf{0.148}\pm 0.097$
($38\pm 1.1$) ($50\pm 1.3$) ($60\pm 1.4$)
$\text{ICt}(1)$ $9.6\cdot 10^{-4}$ $0.101\pm 0.003$ $0.128\pm 0.003$ $0.150\pm 0.003$
($44\pm 1.3$) ($56\pm 1.3$) ($66\pm 1.4$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $2.4\cdot 10^{-3}$ $\textbf{0.085}\pm 0.077$ $\textbf{0.107}\pm 0.077$ $\textbf{0.128}\pm 0.078$
($27\pm 0.8$) ($35\pm 0.9$) ($43\pm 1.0$)
$128\times 128$ $\text{IC}(0)$ $4.7\cdot 10^{-4}$ $1.086\pm 0.163$ $1.367\pm 0.203$ $1.601\pm 0.237$
($147\pm 4.3$) ($185\pm 4.0$) ($217\pm 4.0$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.498}\pm 0.113$ $\textbf{0.628}\pm 0.123$ $\textbf{0.748}\pm 0.134$
($62\pm 2.7$) ($78\pm 2.9$) ($94\pm 3.4$)
$\text{ICt}(1)$ $3.4\cdot 10^{-3}$ $0.629\pm 0.084$ $0.791\pm 0.107$ $0.927\pm 0.125$
($89\pm 2.6$) ($112\pm 2.5$) ($132\pm 2.5$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $5.0\cdot 10^{-3}$ $\textbf{0.344}\pm 0.083$ $\textbf{0.437}\pm 0.093$ $\textbf{0.525}\pm 0.101$
($43\pm 1.5$) ($55\pm 1.8$) ($66\pm 1.8$)
A.2 Additional experiments with ICt(5) preconditioner
Table 9: Comparison on diffusion equation with variance $0.1$: classical algorithms for ICt($5$) and PreCor$\big{[}\text{ICt}(1)\big{]}$. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{ICt}(5)$ $1.9\cdot 10^{-3}$ $\textbf{0.070}\pm 0.003$ $\textbf{0.083}\pm 0.003$ $\textbf{0.098}\pm 0.003$
($20\pm 0.6$) ($27\pm 0.4$) ($32\pm 0.4$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $2.3\cdot 10^{-3}$ $0.088\pm 0.094$ $0.112\pm 0.094$ $0.134\pm 0.094$
($23\pm 0.5$) ($31\pm 0.4$) ($38\pm 0.4$)
$128\times 128$ $\text{ICt}(5)$ $7.9\cdot 10^{-3}$ $0.488\pm 0.074$ $0.625\pm 0.095$ $0.736\pm 0.112$
($42\pm 1.1$) ($54\pm 0.8$) ($64\pm 0.7$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $5.0\cdot 10^{-3}$ $\textbf{0.367}\pm 0.069$ $\textbf{0.469}\pm 0.071$ $\textbf{0.560}\pm 0.074$
($37\pm 0.9$) ($48\pm 0.8$) ($58\pm 0.8$)
Table 10: Comparison on diffusion equation with variance $0.5$: classical algorithms for ICt($5$) and PreCor$\big{[}\text{ICt}(1)\big{]}$. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{ICt}(5)$ $2.2\cdot 10^{-3}$ $\textbf{0.068}\pm 0.002$ $\textbf{0.086}\pm 0.002$ $\textbf{0.112}\pm 0.002$
($21\pm 0.8$) ($28\pm 0.7$) ($33\pm 0.7$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $2.4\cdot 10^{-3}$ $0.085\pm 0.077$ $0.107\pm 0.077$ $0.128\pm 0.078$
($27\pm 0.8$) ($35\pm 0.9$) ($43\pm 1.0$)
$128\times 128$ $\text{ICt}(5)$ $8.1\cdot 10^{-3}$ $0.465\pm 0.050$ $0.588\pm 0.062$ $0.689\pm 0.073$
($45\pm 1.4$) ($57\pm 1.2$) ($68\pm 1.2$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $5.0\cdot 10^{-3}$ $\textbf{0.344}\pm 0.083$ $\textbf{0.437}\pm 0.093$ $\textbf{0.525}\pm 0.101$
($43\pm 1.5$) ($55\pm 1.8$) ($66\pm 1.8$)
A.3 Distribution of eigenvalues
A.4 Generalization
[…]
A.5 Scalability
A.6 Dataset description
@@
Grid $32\times 32$ Grid $64\times 64$ Grid $128\times 128$
@@
$A$ $1024$ 0.4761 $4096$ 0.1205 $16384$ 0.0303
$L$ from IC(00) 0.2869 0.0725 0.0182
$L$ from ICt($1$) 0.3785 0.0961 0.0242
$L$ from ICt($5$) 0.7547 0.1920 0.0485
A.7 Incomplete LU factorization
$A\approx LU,$ (11)
where $LU-A$ satisfies certain constraints.
The complexity of solving sparse linear systems with matrices in the form of the Choletsky decomposition defined by the number of non-zero elements $\mathcal{O}(\text{nnz})$. This value also defines the storage complexity and the complexity of preconditioner construction.
A.8 Details on correction coefficient
Table 13: Values of the learned coefficient $\alpha$ from (6) with loss (3).
Grid $32\times 32$ Grid $64\times 64$ Grid $128\times 128$
Equation Variance IC(0) ICt(1) IC(0) ICt(1) IC(0) ICt(1)
Poisson $-0.159$ $-0.091$ $-0.188$ $-0.095$ $-0.202$ $-0.116$
Diffusion $0.1$ $-0.107$ $-0.073$ $-0.099$ $-0.082$ $-0.115$ $-0.086$
Diffusion $0.5$ $-0.068$ $-0.034$ $-0.064$ $-0.051$ $-0.094$ $-0.055$
Diffusion $0.7$ $-0.066$ $-0.031$ $-0.070$ $-0.043$ $-0.062$ $-0.031$

### v3

@@ 5.1 Preconditioners comparison
We observe a good generalization of our approach when transferring our preconditioner between grids and datasets (see Appendix A.10). The transferability of the PreCorrector allows both to use the pre-trained PreCorrector on similar problems and to train the PreCorrector on the easier problems with inference on the hard ones.
Figure 2: Ablation study of the PreCorrector’s architecture. Top row - preconditioners are constructed from IC(00). Bottom row - preconditioners are constructed from ICt($1$). PreCor w/ GNN - PreCorrector architecture described in Section 3. PreCor w/ MLP - PreCorrector processor block is changed to MLP. PreCor w/ MLP, static diag - same as before without updating the main diagonal. Lower is better.
@@
Results for GNN from (Li et al., 2023) are obtained with the same GNN hyperparameters and architecture as reported in the original paper with only one change: GNN is initialized with all weights and biases equal to zero and trained untill convergence.
Without this change in architecture, the training did not converge. We also observed unstable training of the GNN from (Li et al., 2023) and unpredictable quality of the preconditioner depending on the training length. Incorporating of a normalization used in PreCorrector did not improve the quality of the GNN. In our experiments, we could get a better preconditioner if we did not train until convergence. However, we could not predict the time for early stopping except by directly computing $\kappa(P^{-1}A)$, which is too expensive.
5.2 Ablation study
Learning to correct ILU factors and using a better loss function are key components that allow PreCorrector to produce better preconditioners than classical ones. To further support this claim, we perform an ablation study of these components.
Processor layer of GNN
We have changed the message-passing processor block of PreCorrector (see Section 3) to a simple MLP, resulting in a feed-forward network without any propagation of neighbourhood information (Figure 2).
Experiments on the IC(00) preconditioner show that PreCorrector without message-passing generally produces weaker preconditioners than the architecture described in Section 3 and degradates with increasing coefficient contrast. Nevertheless, it achieves the desired generalization on unseen linear systems and outperforms classical IC(00). On the other hand, the PreCorrector without message-passing cannot handle varying sparsity patterns (experiments with ICt($1$), bottom row, Figure 2) when such preconditioners are given as input matrix $L$.
By pinning diagonal elements $L_{ii}$ in PreCorrector without message-passing, one can stabilize the learning with varying coefficients at the cost of further performance degradation. Diagonal pinning during training can be seen as a variation of the SPD constraint on the matrix $L(\theta)$ that was first introduced in (Sappl et al., 2019) and further used in (Li et al., 2023; Häusner et al., 2023).
The classical ILU algorithms are non-local by design, which allows PreCorrector to construct better preconditioner than their classical analogues for certain problems even with shallow two-layer MLP with $50$ parameters as the processor supporting the efficiency of the proposed approach.
@@ 5.2 Ablation study
Loss function
As mentioned in Section 2, one should focus on approximating the low frequency components. The loss (3) does indeed reduce the distance between the extreme eigenvalues compared to IC(0) which can be observed in Table 5. Moreover, the gap between the extreme eigenvalues is covered by the increase in the minimum eigenvalue, which supports the hypothesis of low frequency cancellation. The maximum eigenvalue also increase, but by a much smaller order of magnitude. In Table 5 we have also included recently obtained bounds for the minimum and maximum eigenvalues (Häusner et al., 2024).
At the same time, the preconditioner trained with the loss function (2) without weighting with $A^{-1}$ gives a worse effect on spectrum of $A$ (Table 6) and takes much more time to converge (Appendix A.8). Different distributions of the eigenvalues can be found in Figure 11 in the Appendix A.9
@@ 6 Discussion
We observe that training a GNN from scratch can be unstable, resulting in preconditioners that have a weaker effect on the spectrum than their classical analogues. In addition, as the matrix size increases, the very first step of training, when GNN is initialized with random weights, leads to an loss overflow since the residual with random $P$ is huge. Learning corrections to classical methods mitigates these problems.
@@ 7 Limitations
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
A.1 Incomplete LU factorization
$A\approx LU,$ (10)
where $LU-A$ satisfies certain constraints.
The complexity of solving sparse linear systems with matrices in the form of the Cholesky decomposition defined by the number of non-zero elements $\mathcal{O}(\text{nnz})$. This value also defines the storage complexity and the complexity of preconditioner construction.
A.2 PreCorrector’s architecture
Algorithm 1 PreCorrector’s forward pass
  Input: IC(00) factor $L$ of $A$, where $e_{ij}^{0}=L_{ij}$
  $\text{s}=\max{(|e_{ij}^{0}|)}$
  $e_{ij}^{1}=\gamma_{\theta}(e_{ij}^{0}~{}/~{}\text{s})$
  $h_{i}^{1}=\textbf{1}$
  for $t=1$ to $T$ do
     $h_{i}^{t+1}=\bigoplus_{j\in\mathcal{N}(i)}h_{j}^{t}~{}e_{ij}^{t}$
     $e_{ij}^{t+1}=\phi_{\theta}\big{(}e_{ij}^{t},h_{i}^{t+1},h_{j}^{t+1}\big{)}$
  end for
  $e_{ij}^{T+1}=\text{s}\cdot\psi_{\theta}(e_{ij}^{T+1})$
  $e_{ij}^{T+1}=e_{ij}^{0}+\alpha\cdot e_{ij}^{T+1}$
  Return: $L(\theta)$, where $L(\theta)_{ij}=e_{ij}^{T+1}$
By using 1 as an input nodes to the processor block, we discard the node encoder MLP. We also notice that node model in processor does not contribute when there are no explicit nodes. We discard it either to use simple aggregation as the node update model and to further reduce number of parameters. Moreover, we make our approach consistent with classical methods by passing only the left-side matrix $A$ to the preconditioner construction routine. We use the max aggregation function, although our experiments showed that the resulting preconditioner quality does not depend on the type of aggregation function.
We also investigate which protocol to follow to increase the receptive field. It is possible to follow the original message passing paper (Gilmer et al., 2017) and perform $T$ rounds of message passing with shared learnable functions during each round. Another option is to follow classic GNN’s multi-block protocol and combine multiple message-passing layers to increase the receptive field with independent weights. During the PreCorrecotr ablation study we observed no difference in resulting preconditioner between the protocols, so we use message-passing with shared layer to further simplify the model.
Table 7: Number of learning parameters for different neural preconditioner routines. PreCor - PreCorrector’s architecture presented in Algorithm 1. PreCor w/ nodes - PreCorrector with node encoder and with node MLP model in Processor. PreCor MB - PreCorrector with multi-block message-passing protocol. PreCor w/o MP - PreCorrector without message-passing (with single MLP).
PreCor PreCor w/ nodes PreCor MB PreCor w/o MP
$1170$ $2753$ $3474$ $50$
A.3 Dataset description
For our experiments we use a following list of computational grids: $\{32,~{}64,~{}128,~{}256\}$. The contrast in the diffusion equation is controlled with a variance in the coefficient function GRF and takes value from $\{0.1,~{}0.5,~{}0.7,~{}1.1\}$. The Gaussian covariance model is used in GRF. The forcing term $f$ is sampled from the standard normal distribution $\mathcal{N}(0,~{}1)$ and each PDE is discretized using the 5-point finite difference method.
@@ A.3 Dataset description
Grid $32\times 32$ Grid $64\times 64$ Grid $128\times 128$ Grid $256\times 256$
@@
$A$ $1024$ 0.4761 $4096$ $0.1205$ $16384$ $0.0303$ $65536$ $0.0076$
$L$ from IC(00) $0.2869$ $0.0725$ $0.0182$ $0.0046$
$L$ from ICt($1$) $0.3785$ $0.0961$ $0.0242$ $0.0061$
$L$ from ICt($5$) $0.7547$ $0.1920$ $0.0485$ $0.0121$
Table 10: Condition number of a sampled linear systems from the used datasets.
Grid
$64$ $128$ $256$
Variance $0.5$ $9.9\cdot 10^{3}$ $5.8\cdot 10^{4}$ $3.5\cdot 10^{5}$
$0.7$ $2.7\cdot 10^{4}$ $8.8\cdot 10^{4}$ $6.0\cdot 10^{5}$
$1.1$ $42\cdot 10^{4}$ $1.8\cdot 10^{5}$ $6.4\cdot 10^{5}$
A.4 Experiments environment
Each dataset from the Section 4 consists of $1000$ training and $200$ test linear systems. The final neural networks are trained with batch size $8$, learning rate $10^{-3}$ and Adam optimizer. Final GNN architecture consists of $5$ message passing rounds and $2$ hidden layers with $16$ hidden features in all MLPs in each experiment. PreCorrector training always starts with the parameter $\alpha=0$ in (6). For training, we used libraries from the JAX ecosystem: jax (Bradbury et al., 2018), optax (DeepMind et al., 2020), equinox (Kidger & Garcia, 2021). We used a single GPU Nvidia A40 48Gb for training. The construction time of preconditioners with neural design was measured on the same GPU. Preconditioners with classical algorithms were generated on the Intel(R) Xeon(R) Gold 6342 CPU @ 2.80GHz with ilupp library [^2] https://github.com/c-f-h/ilupp (Mayer, 2007). The CG method was run on the same CPU using the scipy (Virtanen et al., 2020) implementation.
A.5 Scalability
A.6 Additional experiments with IC(0) and ICt(1) preconditioners
Figure 4: Residuals vs CG iterations for grid $64\times 64$. Lower is better.
Figure 5: Wall time vs CG iterations for grid $64\times 64$. Lower is better.
Figure 6: Residuals vs CG iterations for grid $128\times 128$. Lower is better.
Figure 7: Wall time vs CG iterations for grid $128\times 128$. Lower is better.
Figure 8: Residuals vs CG iterations for grid $256\times 256$. Lower is better.
Figure 9: Wall time vs CG iterations for grid $256\times 256$. Lower is better.
A.7 Additional experiments with ICt(5) preconditioner
Table 11: Comparison on diffusion equation with variance $0.5$: classical algorithm ICt($5$) and PreCor$\big{[}\text{ICt}(1)\big{]}$. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$ to $10^{-12}$
$64\times 64$ $\text{ICt}(5)$ $2.1\cdot 10^{-3}\pm 2.2\cdot 10^{-4}$ $\textbf{0.078}\pm 0.003$ $\textbf{0.099}\pm 0.003$ $\textbf{0.118}\pm 0.003$ $\textbf{0.134}\pm 0.003$
($22\pm 0.7$) ($28\pm 0.6$) ($34\pm 0.6$) ($39\pm 0.7$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $1.0\cdot 10^{-3}\pm 2.6\cdot 10^{-5}$ $0.080\pm 0.008$ $0.103\pm 0.010$ $0.124\pm 0.012$ $0.145\pm 0.013$
($29\pm 1.1$) ($37\pm 1.2$) ($45\pm 1.2$) ($53\pm 1.2$)
$128\times 128$ $\text{ICt}(5)$ $7.9\cdot 10^{-3}\pm 5.6\cdot 10^{-4}$ $0.582\pm 0.066$ $0.737\pm 0.082$ $0.866\pm 0.096$ $0.977\pm 0.108$
($46\pm 1.4$) ($59\pm 1.4$) ($69\pm 1.3$) ($78\pm 1.3$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $3.6\cdot 10^{-3}\pm 1.9\cdot 10^{-4}$ $\textbf{0.432}\pm 0.040$ $\textbf{0.552}\pm 0.051$ $\textbf{0.700}\pm 0.060$ $\textbf{0.788}\pm 0.070$
($47\pm 1.8$) ($60\pm 2.1$) ($73\pm 2.3$) ($85\pm 2.6$)
$256\times 256$ $\text{ICt}(5)$ $3.0\cdot 10^{-2}\pm 1.6\cdot 10^{-3}$ $4.776\pm 0.410$ $5.958\pm 0.507$ $6.942\pm 0.583$ $7.817\pm 0.650$
($97\pm 2.7$) ($121\pm 2.9$) ($142\pm 2.8$) ($160\pm 2.6$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $1.4\cdot 10^{-2}\pm 6.9\cdot 10^{-4}$ $\textbf{2.043}\pm 0.203$ $\textbf{2.567}\pm 0.241$ $\textbf{3.053}\pm 0.273$ $\textbf{3.519}\pm 0.303$
($82\pm 4.5$) ($103\pm 5.2$) ($122\pm 5.7$) ($141\pm 6.2$)
Table 12: Comparison on diffusion equation with variance $0.7$: classical algorithm ICt($5$) and PreCor$\big{[}\text{ICt}(1)\big{]}$. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$ to $10^{-12}$
$64\times 64$ $\text{ICt}(5)$ $2.1\cdot 10^{-3}\pm 1.9\cdot 10^{-4}$ $\textbf{0.080}\pm 0.003$ $\textbf{0.102}\pm 0.003$ $\textbf{0.120}\pm 0.003$ $\textbf{0.136}\pm 0.003$
($23\pm 0.8$) ($29\pm 0.9$) ($34\pm 0.8$) ($39\pm 0.8$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $1.4\cdot 10^{-3}\pm 2.7\cdot 10^{-5}$ $0.075\pm 0.003$ $0.096\pm 0.004$ $0.116\pm 0.004$ $0.135\pm 0.004$
($31\pm 1.4$) ($40\pm 1.6$) ($48\pm 1.7$) ($56\pm 1.7$)
$128\times 128$ $\text{ICt}(5)$ $7.5\cdot 10^{-3}\pm 4.2\cdot 10^{-4}$ $0.661\pm 0.053$ $0.829\pm 0.065$ $0.968\pm 0.075$ $1.088\pm 0.084$
($49\pm 1.7$) ($61\pm 1.7$) ($71\pm 1.8$) ($80\pm 1.7$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $3.6\cdot 10^{-3}\pm 1.9\cdot 10^{-4}$ $\textbf{0.470}\pm 0.056$ $\textbf{0.591}\pm 0.070$ $\textbf{0.710}\pm 0.084$ $\textbf{0.823}\pm 0.097$
($54\pm 2.9$) ($69\pm 3.4$) ($83\pm 3.8$) ($96\pm 4.2$)
$256\times 256$ $\text{ICt}(5)$ $3.0\cdot 10^{-2}\pm 1.6\cdot 10^{-3}$ $4.800\pm 0.624$ $5.980\pm 0.764$ $6.938\pm 0.874$ $7.776\pm 0.968$
($102\pm 3.4$) ($127\pm 3.4$) ($147\pm 3.3$) ($165\pm 3.3$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $1.4\cdot 10^{-2}\pm 6.8\cdot 10^{-4}$ $\textbf{2.800}\pm 0.357$ $\textbf{3.491}\pm 0.434$ $\textbf{4.126}\pm 0.509$ $\textbf{4.735}\pm 0.576$
($90\pm 7.0$) ($113\pm 8.3$) ($134\pm 9.4$) ($154\pm 10.4$)
A.8 Losses comparison
[…]
A.9 Distribution of eigenvalues
A.10 Generalization
[…]